TMS Meeting Minutes, December 10, 2020

Attendees: Andrew, Heejoo, John, Trenton, Matthieu, Olga, Yang (Zoom Meeting)
  • LBT has two dormant laser trackers
  • Andrew thinking that it would be beneficial to have some active laser trackers to do backup measurements
  • Floating this idea, suggests that folks think about it for next meeting

  • Software updates:
  • Matthieu updated the 'tms' keyword on DMS
  • Now able to look up laser power, etc. with ease
  • Thank you, software team!
  • Also, softare team working on an email alert for laser leak
  • The alarm should be ready in 1-2 weeks

  • Laser now on continuously, since Dec. 4
  • Heejoo monitoring the laser power and has found now leakage
  • Planning to keep the laser on until next week
  • If this doesn't pose an issue, Heejoo proposes that the TMS laser remain on until the laser might affect observation (LUCI h-band, LBTI)
  • Andrew suggests daytime commissioning activities where we run long exposures with TMS laser on to determine if there are any stray light contributions
  • So far, a 10 minute exposure on LUCI h-band can not see the lasers, however these tests were conducted without the LUCI configuration setup (M1 to M2)
  • Olga suggests that testing this stray light with the M1-M2-M3-LUCI system would be a good task for a night with poor weather, perhaps tonight
  • Start with alignment to open the shutter, take exposure
  • Lasers on shutters closed, take exposure
  • Lasers off, take exposure
  • Expect to see indistinguishable results from lasers off to laser on-closed shutter
  • Olga planning to support this test
  • LUCI, hk-spectrum, the laser shows up as two distinct lines
  • John thinks there is not much difference between imaging and spectra
  • Imaging, more photons, but they are more spread out
  • Spectra, fewer photons, but the analysis of the result is very clear
  • Heejoo and Olga planning to start after 9:00 PM tonight

  • Heejoo is working on data processing for correlation of day-by-day pose of LBC
  • Gathering LBC images and comparing raw data from TMS with temperature, rel. humidity
  • Attempting to find correlation between best IQ and pose data/collimation
  • Hoping to obtain absolute pose (against the first day, permanent reference pose), elevation, IQ correlation
  • Heejoo setting first day pose as the reference
  • Andrew: must be very careful to get the Rx, Ry terms which are deliberately injected subtracted from the data; would also like to see 5-shot running average of mirror diameter
  • Want to see if there is a correlation of glass global scale and focus

  • Trenton working on comparison plot between IQ, TMS data, and telemetry
  • Still working on this
  • Elevation instead of azimuth suggested
  • Things desired on plot:
    • Mode of software operation (dropped vs. no dropped channels)
    • Mirror diameter (unfortunately, not collected with the current TMS code)
    • Request for new plot with the time-derivative of the temperature telemetry
    • Filter out to only include g-SLOAN and i-SLOAN, so as not to include the effect of wavelength change
    • Cumulative FPIA correction (astigmatism, spherical aberration, and focus); PSF channel on DMS should include FPIA zernikes

  • Issues with the TEC chip focus
  • TEC chip noise issues have resulted in problems
  • There is a gradient of FWHM across the detector; donuts at the top of chip 2, elongated at the bottom
  • Notable donuts when ch. 17 was dropped
  • The science night during which this effect was seen, the temperature was dropping
  • John has been working on the jumps, and he believes there might be a bug in PSF
  • The group rotation correction is fighting against the TMS corrections (only happens when TMS is running)
  • The particular jumps at 9:40, 10:12, 12:40 are likely caused by PSF bug

  • Andrew has been speaking with Joel Nyssen about the need or lack thereof for correction vectors
  • Joel Nyssen is surprised that we need a correction vector
  • Believes something funny is going on
  • Andrew working through the underlying assumptions for using the dropped channel correction vector

  • SPIE Astronomical Instrumentation starts next week
  • Heejoo and Olga will run a laser leakage test with LUCI tonight (spectral, imaging)
  • Note, not typical to take a 10 minute exposure with LUCI (Dave tested this case as an edge case to prove there is no stray light)
  • TeamViewer issues
  • Currently, tms_extract does not extract the pose data
  • Yang adding this functionality to the tms_extract utility

-- TrentonBrendel - 10 Dec 2020
Topic revision: r2 - 04 Jan 2021, TrentonBrendel
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback